# Channel-Level Calibration Methods of Silicon Photomultiplier for JUNO-TAO Central Detector

### v1

@@ Abstract
The Taishan Antineutrino Observatory (TAO or JUNO-TAO) is a satellite observatory for the Jiangmen Underground Neutrino Observatory (JUNO), located 44 meters away from the No.1 reactor of the Taishan Nuclear Power Plant. TAO can measure the reactor antineutrino energy spectrum with excellent energy resolution (better than 2% at 1 MeV) using state-of-the-art Silicon Photomultipliers (SiPMs) operated at low temperature. To achieve this goal, the SiPMs (together with their readout electronics) must be well calibrated. This paper presents the channel-level calibration methods for the dark count rate (DCR), relative photon detection efficiency (PDE), time offset, gain, and internal optical crosstalk (IOCT) of the SiPMs based on charge and time information of the collected events. For the calibration of the external optical crosstalk (EOCT), in terms of its rate and emission angle distribution, a novel method is proposed by switching on and off different groups of SiPMs with an LED placed in the detector. Using one million simulated events, the expected calibration biases are evaluated for all the aforementioned parameters: relative PDE (2.1%), IOCT (1.4%), DCR (0.4%), EOCT Rate ($0.1\%$), gain ($0.1\%$), time offset (0.027 ns). The emission angle distribution of the EOCT photons could be measured with a bias of less than 4% in main angular range.
@@ 1 Introduction
Before the precise measurements of the reactor neutrino energy spectrum by experiments such as Daya Bay [Ref1], RENO [Ref2], and Double Chooz [Ref3], the primary theoretical models for calculating the expected reactor neutrino energy spectrum were the Summation model [Ref4] based on nuclear database data [Ref33, Ref34, Ref35, Ref36, Ref37, Ref38] and the $\beta$ conversion method such as Huber-Muller model [Ref5, Ref6] based on the ILL High Flux Reactor measurement data [Ref7, Ref8, Ref9]. However, according to the measurement results from the Daya Bay, RENO, and Double Chooz experiments, both the Summation model and the Huber-Muller model exhibit systematic deviations from the experimental results in terms of the predicted energy spectrum flux and the shape of the energy spectrum at 5 MeV. Although Estienne and Fallot have improved the Summation model [Ref10]; and Kopeikin, Skorokhvatov, Titov [Ref11]; Hayen, Kostensalo, Severijns, Suhonen [Ref12]; Giunti, Li, Ternes, Xin [Ref13] have refined the Huber-Muller model, these models currently exists that can not simultaneously explain the systematic deviations in both flux and spectrum shape. Moreover, none of models can reliably predict the fine structures in the reactor antineutrino spectrum [Ref46]. If the JUNO experiment directly uses these predicted spectra as inputs, it will introduce significant uncertainties in the energy spectrum, which will affect the sensitivity of the neutrino mass ordering measurement [Ref14]. To reduce the model dependence of the reactor neutrino energy spectrum, JUNO has deployed a satellite detector, the Taishan Antineutrino Observatory (TAO or JUNO-TAO) at a baseline of 44 meters away from Unit 1 of the Taishan Nuclear Power Plant [Ref15].
@@
To achieve the aforementioned physical goals, it is crucial to attain the designed energy resolution for the TAO experiment. However, the negative effects of SiPMs can directly impact the energy resolution. For instance, according to the TAO Conceptual Design Report, DN and optical crosstalk (OCT) contribute to a degradation of 0.75% and 0.5% on the energy resolution at 1 MeV [Ref15]. In addition to affecting the energy resolution, these two effects can cause deviations of 1 mm and 2 mm on the vertex reconstruction, as reported by Shi [Ref16]. Therefore, accurate calibration of all relevant SiPM parameters for each readout channel is of crucial importance.
@@
This paper investigates and presents channel-level calibration methods for SiPMs installed in the TAO CD, including calibration of the dark count rate (DCR), relative photon detection efficiency (PDE), internal optical crosstalk (IOCT) rate, gain, and time offset. Additionally, a method based on selectively turning on and off different groups of SiPMs utilizing the fine control capability of the SiPM high-voltage system and an LED calibration source is proposed to calibrate the EOCT rate and emission angle distribution. This study is conducted using simulated data generated by the TAO offline software.
This paper is structured as follows: Section 2 introduces the TAO detector; Section 3 presents the simulated model; Section 4 describes the calibration of the DCR, relative PDE, and time offset based on hit time information; Section 5 details the calibration of gain and IOCT rate based on charge information; Section 6 introduces a novel method for EOCT rate and emission angle distribution calibration; Section 7 summarizes the results and discusses limitations of these methods.
@@ 2 TAO Detector
Figure 1: Design of the TAO detector
The central detector contains 2.8 tonnes of GdLS [Ref20] enclosed within an acrylic sphere. The exterior of the acrylic sphere is covered by a copper shell, which inner surface is equipped with 4024 HPK SiPMs, each measuring $50.7\times 50.7$ ${\mathrm{mm}}^{2}$. The copper shell is housed inside a stainless steel tank filled with linear alkylbenzene, which serves as a buffer liquid to shield external radioactive background, maintain temperature stability, and provide optical coupling between the acrylic container and the SiPMs. Both the tank and the copper shell have refrigeration tubes that cool the central detector to -50 °C to reduce the SiPM DCR to below 100 ${\mathrm{mm}}^{2}$. At the top of the CD is the calibration system [Ref21], which includes an Automatic Calibration Unit (ACU) and a Cable Loop System (CLS). The ACU can deploy calibration sources (include LED source, $^{68}{\rm Ge}$ source and combined $\gamma$ sources) along the central axis of the GdLS. It is used to calibrate the CD’s non-linearity and perform channel-level calibration of SiPMs. The CLS consists of the stainless steel cable passing through specific anchor points with a small segment plated with $^{137}{\rm Cs}$, which is used to calibrate the non-uniformity of the central detector.
@@
Above the central detector is the top veto tracker [Ref47, Ref48], which consists of four layers of plastic scintillators and is designed for muon veto.
@@ 3 TAO Central Detector Simulated Model
This study utilizes simulated data generated by the TAO offline software to investigate calibration methods for SiPM parameters. This section presents the simulation model.
TAO offline software is a tool for experimental data processing, physics data analysis and simulated data production. As a key component of the TAO experiment, the TAO offline software is built upon the SNiPER [Ref23] framework and includes modules for physics event generator, detector simulation, electronics simulation, raw data conversion, calibration, reconstruction, and physical analysis. For simulated data production, the physics event generator, detector simulation and electronics simulation modules are utilized.
@@ 3.1 Physics Event Generator
The physics event generator is used to produce particles and specific events, and the physics event generator of TAO directly adopts that of JUNO. In JUNO simulation paper, it indicates that we have three types physics event generators: particle gun, HepEvt and GENIE [Ref24]. In this paper, we use particle gun which is a tool used to generate particles with specific particle types, momenta and positions.
@@ 3.2 Detector Simulation
Following the physics event generator, the detector simulation is performed. The TAO detector simulation is based on Geant4 (version:10.04.p02) [Ref41], and models the propagation of particles through the detector. The TAO detector simulation module is largely derived from the Daya Bay detector simulation and has been modified in the following four aspects:
  1. 1.
    Detector geometry construction: The detector geometry has been changed to TAO’s [Ref15].
  2. 2.
    Physics processes: The physics process has been changed to JUNO’s [Ref24].
  3. 3.
    Optical model of the LS: The optical model is constructed based on three components: linear alkylbenzene (LAB), 2,5-diphenyloxazole (PPO) and p-bis-(o-methylstyryl)-benzene (bis-MSB) [Ref49]. The absorption-remission spectrum is determined by PPO and bis-MSB [Ref50], and the light attenuation length is jointly determined by the absorption probabilities of LAB, PPO, and bis-MSB. Additionally, the Birks constant has been updated to match that used in the JUNO simulation.
  4. 4.
    Sensitive detector: The sensitive detector has been changed from PMTs to SiPMs. The photon detection efficiency and reflectivity have been configured according to experimental measurements, where the photon detection efficiency is about 50% and the reflectivity is about 8.5%. [Ref32, Ref51].
@@ 3.3 Electronics Simulation
The electronics simulation module is designed to simulate the process from SiPM photoelectrons to the readout of electrical signals by the trigger system. It primarily consists of three components: SiPM effect modeling, event triggering, and waveform simulation.
3.3.1 SiPM Effect Modeling
When events from the detector simulation are input into the electronics simulation module, each hit in each event is converted into an electronic pulse. The electronics simulation module add SiPM effects, which including dark noise (DN), gain, time offset, optical crosstalk (OCT), and afterpulse (AP).
@@ 3.3.1 SiPM Effect Modeling
The gain determines the amplitude of each pulse. The amplitude of each pulse is sampled from a Gaussian distribution with a mean value of 1 and the standard deviation of 15% corresponding to the charge resolution of electronics system [Ref15]. The amplitude of the pulse indicates how many times greater the waveform amplitude is that of a single photoelectron (SPE) waveform. By integrating the SPE waveform, the true gain value can be determined which is 3000 ADC counts.
Regarding the time offset, the timestamp of each pulse is adjusted by adding a constant time offset specific to the corresponding channel. In this study, a 20 ns time offset is applied to the SiPMs located in the northern hemisphere. Additionally, a Gaussian smearing with a standard deviation of 1 ns [Ref15] is added to each pulse timestamp for time spread.
For the OCT, each pulse can generate OCT pulses (including the OCT itself), and the number of OCT pulses generated by each pulse is determined via Poisson sampling. OCT is divided into internal optical crosstalk (IOCT) and external optical crosstalk (EOCT). The IOCT refers to OCT pulses occurring in the same channel as the source pulse, while the EOCT occurs in different channels. In the TAO electronics simulation, IOCT and EOCT sampling are independent, each controlled by separate Poisson parameters, and every pulse undergoes both IOCT and EOCT sampling. During the TAO SiPM mass test, the SiPMs were tested in air; however, the OCT rates in air and LAB are different. Therefore, the OCT parameters used in the simulation are not from the TAO SiPM mass test, they are from Guan’s paper, which measured the SiPM OCT rate in the LAB [Ref17]. However, in his paper, the IOCT rate is very low, making it unsuitable for studying the IOCT calibration method. Consequently, we set the IOCT Poisson parameter to 0.15. The EOCT Poisson value is 0.2447 multiplied by the detector geometric factor, which corrects for the fact that the SiPM coverage is not 100%. The detector geometric factors vary depending on each channel, with values ranging from 0.7 to 0.8.
For the AP, each main pulse may generate an AP. The occurrence of an AP is determined by sampling from a Bernoulli distribution with an AP rate of 1%, as obtained from the TAO SiPM mass test [Ref32]. The delay of the AP relative to the main pulse follows an exponential distribution with a parameter of 50 ns. The delay time interval between the AP and the main pulse determines the amplitude of the AP, which is given by the expression $(1-e^{\frac{-\Delta t}{\tau}})$, where $\Delta t$ is the time interval and $\tau$ is the recovery time of the SiPM, set to 32.5 ns. These two time constants are from Rolph’s paper [Ref52].
The parameters are summarized in Table 1. The typical ranges for DCR, PDE and AP are based on the TAO SiPM mass test result [Ref32]. The values for charge resolution and time spread are taken from our conceptual design report [Ref15]. The typical ranges for internal and external OCT are sourced from Guan’s paper [Ref17]. The gain value is derived from our prototype detector result [Ref53].
Parameter Typical Range Simulated

### v2

@@ Abstract
The Taishan Antineutrino Observatory (TAO or JUNO-TAO) is a satellite observatory for the Jiangmen Underground Neutrino Observatory (JUNO), located 44 meters away from the No.1 reactor of the Taishan Nuclear Power Plant. TAO can measure the reactor antineutrino energy spectrum with excellent energy resolution (better than 2% at 1 MeV) using state-of-the-art Silicon Photomultipliers (SiPMs) operated at low temperature. To achieve this goal, the SiPMs (together with their readout electronics) must be well calibrated. This paper presents the channel-level calibration methods for the dark count rate (DCR), relative photon detection efficiency (PDE), time offset, gain, and internal optical crosstalk (IOCT) of the SiPMs based on charge and time information of the collected events. For the calibration of the external optical crosstalk (EOCT), in terms of its rate and emission angle distribution, a novel method is proposed by switching on and off different groups of SiPMs with an LED placed in the detector. Using one million simulated events, the expected calibration biases are evaluated for all the aforementioned parameters: relative PDE (2.1%), IOCT (1.4%), DCR (0.4%), EOCT Rate ($0.1\%$), gain ($0.1\%$), time offset (<0.2 ns). The emission angle distribution of the EOCT photons could be measured with a bias of less than 4% in main angular range.
@@ 1 Introduction
Before the precise measurements of the reactor neutrino energy spectrum by experiments such as Daya Bay [6], RENO [11], and Double Chooz [1], the primary theoretical models for calculating the expected reactor neutrino energy spectrum were the Summation model [14] based on nuclear database data [17, 5, 36, 32, 40, 31] and the $\beta$ conversion method such as Huber-Muller model [22, 30] based on the ILL High Flux Reactor measurement data [38, 37, 19]. However, according to the measurement results from the Daya Bay, RENO, and Double Chooz experiments, both the Summation model and the Huber-Muller model exhibit systematic deviations from the experimental results in terms of the predicted energy spectrum flux and the shape of the energy spectrum at 5 MeV. Although Estienne and Fallot have improved the Summation model [13]; and Kopeikin, Skorokhvatov, Titov [23]; Hayen, Kostensalo, Severijns, Suhonen [21]; Giunti, Li, Ternes, Xin [16] have refined the Huber-Muller model, these models currently can not simultaneously explain the systematic deviations in both flux and spectrum shape. Moreover, none of models can reliably predict the fine structures in the reactor antineutrino spectrum [12]. If the JUNO experiment directly uses these predicted spectra as inputs, it will introduce significant uncertainties in the energy spectrum, which will affect the sensitivity of the neutrino mass ordering measurement [3]. To reduce the model dependence of the reactor neutrino energy spectrum, JUNO has deployed a satellite detector, the Taishan Antineutrino Observatory (TAO or JUNO-TAO) at a baseline of 44 meters away from Unit 1 of the Taishan Nuclear Power Plant [2].
@@
To achieve the aforementioned physical goals, it is crucial to attain the designed energy resolution for the TAO experiment. However, the negative effects of SiPMs can directly impact the energy resolution. For instance, according to the TAO Conceptual Design Report, dark noise and optical crosstalk (OCT) contribute to a degradation of 0.75% and 0.5% on the energy resolution at 1 MeV [2]. In addition to affecting the energy resolution, these two effects can cause deviations of 1 mm and 2 mm on the vertex reconstruction, as reported by Shi [39]. Therefore, accurate calibration of all relevant SiPM parameters for each readout channel is of crucial importance.
@@
This paper investigates and presents channel-level calibration methods for SiPMs installed in the TAO CD, including calibration of the DCR, relative PDE, internal optical crosstalk (IOCT) rate, gain, and time offset. Additionally, a method based on selectively turning on and off different groups of SiPMs utilizing the fine control capability of the SiPM high-voltage system and an LED calibration source is proposed to calibrate the EOCT rate and emission angle distribution. This paper is conducted using simulated data generated by the TAO offline software.
This paper is structured as follows: Section 2 introduces the TAO detector; Section 3 presents the simulated model; Section 4 describes the calibration of the DCR, relative PDE, and time offset based on hit time information; Section 5 details the calibration of gain and IOCT rate based on charge information; Section 6 introduces a novel method for EOCT rate and emission angle distribution calibration; Section 7 discusses the SiPM parameter uncertainties induced by temperature fluctuation and the impact of uncertainties originating from both temperature fluctuation and the calibration medthon on the reconstructed vertex and energy; Section 8 summarizes the research work presented in this paper.
@@ 2 TAO Detector
Figure 1: Design of the TAO detector.
[…]
@@
Above the central detector is the top veto tracker [27, 28], which consists of four layers of plastic scintillators and is designed for muon veto.
@@ 3 TAO Central Detector Simulated Model
This paper adopts simulated data produced with the TAO offline software to investigate calibration methods for SiPM parameters, and this section details the corresponding simulation model.
The TAO offline software is a dedicated framework for experimental data processing, physics data analysis, and simulated data production for the TAO experiment. As a core offline component of the experiment, it is built on the SNiPER framework [26]. For the simulated data production in this work, three core modules are employed: physics event generation, detector simulation, and electronics simulation.
@@ 3.1 Physics Event Generator
The first step of the full Monte Carlo (MC) simulation chain for the TAO is to produce primary particles using the physics event generator. The generator used in the TAO is directly inherited from the well-validated offline framework of the JUNO experiment. As documented in the JUNO simulation paper [25], the JUNO framework supports three primary types of physics event generators: the particle gun, HepEvt, and GENIE.
In this work, we adopt the particle gun, a dedicated tool that generates particles with well-defined species, momenta, and positions. We generate one million particle events for each SiPM parameter, and the particle events information corresponding to each SiPM parameter is summarized in Table 1.
SiPM Parameter Particle Position Trigger Type
Dark Count Rate Arbitrary1 Arbitrary Self Trigger
Gain Arbitrary Arbitrary Self Trigger
Time Offset LED CD Centre External Trigger
Relative PDE $^{68}\text{Ge}$ CD Centre Self Trigger
Internal Crosstalk Arbitrary Arbitrary Self Trigger
External Crosstalk LED CD Centre External Trigger
  • 1
    We use electrons uniformly distributed within the liquid scintillator with energies following a uniform distribution in the range of 1–10 MeV to represent arbitrary events.
@@ 3.2 Detector Simulation
Following the physics event generation, detector simulation is performed as the next step of the simulation chain. The detector simulation module in the TAO offline software is built based on the Geant4 (version:10.04.p02) [4] toolkit. Firstly, the TAO detector geometry is constructed in Geant4 according to the detector structure of the TAO shown in Figure 1. Subsequently, the physics processes of particles propagating inside the detector are defined. All physics processes are consistent with those implemented in the detector simulation of the JUNO [25], except for two dedicated optical processes: the scintillation light emission induced by energy deposition of particles in the gadolinium-doped liquid scintillator, and the physical process of optical photon detection by SiPMs.
For the simulation of optical processes in the liquid scintillator, a quenching effect occurs when charged particles deposit energy in the liquid scintillator. This quenching effect is described by Birks’ law [7], and the Birks constant of the liquid scintillator adopted in the TAO is determined from dedicated experimental measurements [47]. The remaining deposited energy, after excluding the quenched component, drives the scintillation light emission of the liquid scintillator. The wavelength distribution of the emitted scintillation photons is determined by the 2,5-diphenyloxazole (PPO) and p-bis-(o-methylstyryl)-benzene (bis-MSB), and the wavelength spectrum implemented in the TAO detector simulation is fixed based on the experimentally measured in Ref. [49]. The optical attenuation length of the TAO liquid scintillator is jointly determined by the linear alkylbenzene, PPO and bis-MSB, and the value adopted in the TAO simulation is experimentally determined as specified in Ref. [49]. In addition, the light yield and scintillation time profile of the TAO liquid scintillator are experimentally characterized and determined as described in Ref. [44].
For the simulation of optical processes for SiPMs, when optical photons hit on the surface of a SiPM, they can be detected or reflected by the SiPM. The former is determined by the PDE of the SiPM, while the latter is governed by the optical reflectivity of the SiPM. Both parameters implemented in the TAO detector simulation are derived from dedicated experimental measurements: the PDE is taken from Ref. [45], and the optical reflectivity is adopted from Ref. [43].
Finally, the SiPMs detect both the light produced by energy deposition of particles in the liquid scintillator and the Cherenkov light generated via Cherenkov radiation, and store the corresponding detection data as SiPM hit information.
@@ 3.3 Electronics Simulation
Following the generation of SiPM hit information via detector simulation, electronics simulation is performed as the subsequent stage of the full simulation chain. The TAO electronics simulation workflow is illustrated in Figure 2, which takes the SiPM hit information produced by detector simulation as its input. The hit information is fed into the TaoPdSimAlg algorithm, which converts the SiPM hit information from each detector simulation event into electrical pulse data. Each electrical pulse is characterized by three core parameters: relative amplitude, timestamp, and type identifier. The relative amplitude is determined by the single photoelectron (PE) gain and single PE charge resolution of the SiPM. For the timestamp, the event-level timestamp of each electronics simulation event is obtained by sampling based on the physics event rate; the timestamp of each electrical pulse is then calculated by summing the event-level timestamp, the corresponding SiPM hit time, and the contributions from the electronics time offset and time smearing effects. The type identifier is used to indicate the specific SiPM channel or electronics process from which the electrical pulse originates. Thereafter, additional electrical pulses induced by the intrinsic SiPM effects, including dark noise, OCT, and afterpulse (AP), are generated based on the corresponding dedicated models. Details of the SiPM effect models are elaborated in Section 3.4.
Figure 2: TAO Electronics Simulation Workflow.
Following the generation of all electrical pulses, trigger decision and event packaging are performed within the TaoTrigSimAlg algorithm. The TAO electronic system supports two distinct trigger modes: self-trigger mode and external trigger mode. For the self-trigger mode, trigger decision is implemented based on a 300 ns trigger time window. This time window slides along the time axis, when the number of channels in which electrical pulses are detected within the this time window exceeds the trigger threshold (set to 1000 in the TAO electronics simulation), a valid trigger will be generated, and the start time of the trigger time window is defined as the trigger time. For the external trigger mode, the internal threshold-based trigger decision logic is bypassed, and a forced trigger is generated for each individual detector simulation event. The timestamp sampled from the corresponding detector simulation event is directly adopted as the electronics trigger time. For both trigger modes, a fixed readout window spanning from 100 ns before the trigger time to 900 ns after the trigger time is applied, and all electrical pulses falling within this window are packaged into a single electronic event.
Following the generation of electronics event, the event is fed into the TaoElecSimAlg algorithm for waveform simulation, which converts each electrical pulse into an electrical pulse waveform. The amplitude of the waveform is determined by the relative amplitude of the electrical pulse, and white noise with a signal-to-noise ratio of 10 is added to the waveform. Subsequently, SiPM waveform identification and charge-time calculation are performed. A valid SiPM waveform is identified when three consecutive sampling points exceed the preset threshold (set at 50% of the single PE waveform amplitude). The charge corresponding to the waveform is obtained via waveform integration, and the timing of the waveform is determined by linear fitting of the waveform rising edge, where the intersection of the fitted rising edge and the baseline is defined as the waveform timing. After the charge-time calculation is completed, three parameters of each waveform are obtained: the charge, the time relative to the trigger time, and the associated electronics channel number. Finally, the SiPM parameters of each channel are calibrated based on the timing and charge information.
In the TAO simulation, the true values of each SiPM parameter are known. Accordingly, we quantify the calibration performance using two core metrics: the calibration bias and the standard deviation. The calibration bias is defined as the mean difference between the calibrated value and the corresponding simulated truth, averaged over all channels. The calibration standard deviation is calculated as the standard deviation of the residuals (calibrated value minus the corresponding simulated truth) across all channels.
3.4 SiPM Simulation Modeling
3.4.1 Dark Noise
Dark noise refers to the effect in which a SiPM spontaneously generates avalanche current in the absence of photon hits. It arises from two distinct physical processes: one is thermal current, which is generated when valence band electrons in the SiPM are excited into the conduction band by thermal energy, and trigger an electron avalanche to produce avalanche current during this process; the other is tunneling current, which occurs when valence band electrons in the SiPM have a probability of tunneling into the conduction band via the quantum tunneling effect under the action of an electric field, and subsequently induce avalanche current [33].
@@ 3.4.1 Dark Noise
3.4.2 Time Offset
In the TAO experiment, differences in the length of signal cables for SiPMs at different positions, and potential time asynchrony between the ADCs used in the TAO readout system. These effects lead to relative time offsets between different electronics channels. In the TAO electronics simulation, the time offset is implemented by adding distinct time offset constants to the timestamps of electrical pulses corresponding to different channels. In this study, the time offset is configured to 0 ns for the upper hemisphere channels and 20 ns for the lower hemisphere channels.
3.4.3 Photon Detection Efficiency
The SiPM PDE is defined as the ratio of the number of photons detected by the SiPM to the number of photons incident on the SiPM surface. The SiPM PDE is generally characterized by three factors [20, 48]:
$PDE(\lambda,V)=QE(\lambda)\cdot Prob_{\rm GM}(V)\cdot FF$ (3.1)
where QE is the Quantum Efficiency, which varies with the wavelength of incident light; $Prob_{\rm GM}$ is the probability of triggering an electron avalanche effect, which is related to the internal electric field strength of the SiPM and thus to the operating voltage of the SiPM; and FF (Fill Factor) is a geometric factor used to evaluate the impact of the SiPM geometry on the photon detection efficiency. For the TAO simulation, the detector geometry is constructed from engineering drawings, and the geometric effect is automatically set by Geant4.
The PDE used in the TAO detector simulation is set to 50%, which is generally consistent with the measured results from TAO SiPM mass test [42]. This PDE value is configured on the optical surface of the Geant4 sensitive detector corresponding to the SiPMs in the TAO detector simulation.
3.4.4 Gain
Gain is defined as the total charge of the waveform, which is generated after a photon hits the SiPM, is detected, produces an avalanche current, and is amplified and shaped by the electronics system.
Due to various stochastic processes in the SiPM and the electronics, there is a fluctuation in the amplification factor that converts a single photoelectron into a single photoelectron waveform, which is the charge resolution of the SiPM and the electronics. In the TAO electronics simulation, a 15% Gaussian smearing is added to the relative amplitude of the electrical pulse to represent the 15% charge resolution of the TAO SiPM and electronics reported in the TAO Conceptual Design Report [2].
3.4.5 Optical Crosstalk
OCT of the SiPM arises from the fact that avalanche electrons have a finite probability of emitting photons during acceleration in the electric field of the SiPM. These photons have a certain probability of leaving the Avalanche Photodiode (APD) where they are generated, impinging on other APDs, and triggering new avalanche signals [29, 35].
If the APD hit by the OCT photon and the APD that generates the photon belong to the same electronics channel, this effect is defined as IOCT. If the APD hit by the OCT photon and the APD that generates the photon belong to different electronics channels, this effect is defined as EOCT.
According to the study by Vinogradov [41], the number of OCT events generated by a signal follows a cascaded Poisson distribution. For a single avalanche signal, the total number of the primary signal and the OCT events it induces follows a Borel distribution: